OpenStudy (anonymous):

MEDAL! Can the numbers 24,32,40 be the lengths of the three sides of a right triangle? explain why or why not don't get this, please help!

OpenStudy (anonymous):

@horsegirl27

Mehek (mehek14):

right angle triangles have 1 90 degree angle

Mehek (mehek14):



OpenStudy (anonymous):

yes but this is saying they are the side lengths

Mehek (mehek14):

then you have to use pythagorean theorem

OpenStudy (anonymous):

idk what that is, haven't got to that yet:/

Mehek (mehek14):

\(\Large{a^2 + b^2 = c^2}\) this is the general formula

OpenStudy (anonymous):

Isnt the hypotenuse of the triangle the sum of the two legs? correct me if im wrong

Mehek (mehek14):

use 24 for a 32 for b and 40 for c

Mehek (mehek14):

so: \(\color{blue}{24^2 + 32^2 = 40^2}\)

Mehek (mehek14):

what is \(\color{blue}{24^2}\)?

OpenStudy (anonymous):

since I haven't got to this subject in school I cant use this theorem

OpenStudy (anonymous):

What formulas have you been using in school to work with triangles?